OpenStudy (anonymous):

If a=<-3,2,5> and b=<6,2,3>, calculate the equation of the plane parallel to both a and b containing the point P(-1,9,2)

OpenStudy (turingtest):

to get the equation of a plane you need a point and a vector normal to the plane if \(\vec a\) and \(\vec b\) are both parallel to the plane, then their cross product should give the normal vector for the plane

OpenStudy (turingtest):

so find the normal vector\[\vec n=\vec a\times\vec b\]

OpenStudy (anonymous):

-4i+39j-18k= <-4,39,-18> right?

OpenStudy (turingtest):

I think your j component is 45

OpenStudy (anonymous):

Would the equation be: \[-4(x+1)+39(y-9)-18(z-2)\]

OpenStudy (turingtest):

\[\vec a\times\vec b=\left|\begin{matrix}\hat i&\hat j&\hat k\\-3&2&6\\6&2&3\end{matrix}\right|=6\hat i+36\hat j-6\hat k-12\hat k-12\hat i+9\hat j=\langle-6,45,-18\rangle\]

OpenStudy (anonymous):

You put -3,2,6 instead of -3,2,5

OpenStudy (turingtest):

oh ic

OpenStudy (turingtest):

\[\vec a\times\vec b=\left|\begin{matrix}\hat i&\hat j&\hat k\\-3&2&5\\6&2&3\end{matrix}\right|=6\hat i+30\hat j-6\hat k-12\hat k-10\hat i+9\hat j=\langle-4,39,-18\rangle\]ok you got it, just be sure top set what you wrote equal to zero

OpenStudy (anonymous):

Thanks!

OpenStudy (turingtest):

welcome :)